OpenStudy (anonymous):

What is the x-coordinate of the point on the graph of the equation y=2x+3 which is closest to the point (5,0)?

OpenStudy (anonymous):

The closest distance means the distance square will be minimum as well. so use d^2 = (x - 5)^2 + ((2x+3) - 0)^2

OpenStudy (anonymous):

Then try to minimize this.

OpenStudy (anonymous):

So I used the distance formula r= sqrt((x1-x2)^2+(y1-y2)^2) r=sqrt((x1-5)^2+(y1-0)^2) and followed up using y=2x+3 r=sqrt((x-5)^2+(2x+3)^2)

OpenStudy (anonymous):

By the way, I'm using the calculator method, not calc.

OpenStudy (anonymous):

You don't need calc or even a calculator to do this.

OpenStudy (anonymous):

how else would you find the minimum?

OpenStudy (anonymous):

You can't used -b/2a, can you?

OpenStudy (anonymous):

The definition of a parabola is the distance between some point (focus) and some line (directrix), so we know we have a parabola. A parabola always has its critical point at the vertex.

OpenStudy (anonymous):

Alright, then. How would you minimize this: r=sqrt((x-5)^2+(2x+3)^2)?

OpenStudy (anonymous):

It leads to sqrt(5x^2+2x+34), by the way.

OpenStudy (anonymous):

Sure, well first square both sides. We know that for all cases where r < r', then r^2 < r'^2

OpenStudy (anonymous):

This is because we know r is positive.

OpenStudy (anonymous):

So then we have: r^2 = 5x^2+2x+34 Complete the square, and you get something of the form: r^2 = a(x+h)^2 + k We know the vertex is going to be (-h,k)

OpenStudy (anonymous):

We'll know that -h is the x coordinate of the closest point, and put it into y=2x+3 to get the y coordinate

OpenStudy (anonymous):

but can we ignore the ^2 on the distance when finding the vertex?

OpenStudy (anonymous):

Okay, let me explain this again. Suppose we know that r is the smallest possible distance.

OpenStudy (anonymous):

This means that for all r', we know r < r'

OpenStudy (anonymous):

Suppose we square both sides. Since we know that r and r' are non-negative (for the fact they are distances), we know that r^2 < r' * r < r'^2

OpenStudy (anonymous):

In other worse, squaring will preserve order and the minimum square is still the minimum.

OpenStudy (anonymous):

I just wasn't sure if the x-input value would be preserved.

OpenStudy (anonymous):

Is completing the square too much? Are inequalities too much?

OpenStudy (anonymous):

The x input doesn't need to be preserved.

OpenStudy (anonymous):

The x input that minimizes r will minimize r^2

OpenStudy (anonymous):

I understand the inequalities, though I'm a bit rusty on changing quadratic functions to vertex form.

OpenStudy (anonymous):

That is, I don't know how to do it if the leading coefficient isn't 1. Help? :D

OpenStudy (anonymous):

Divide the whole thing by 5.

OpenStudy (anonymous):

now the leading coefficient is 1. Multiply the 5 back in when you're done.

OpenStudy (anonymous):

Oh wait. Nevermind. Purplemath to the rescue! Thanks for everything.

OpenStudy (anonymous):

Did you write that, or is it a frightful coincidence?

OpenStudy (anonymous):

I wrote it a while ago

OpenStudy (anonymous):

oh cool.

OpenStudy (anonymous):

Like a couple months ago

OpenStudy (anonymous):

Are you planning to finish the website?

OpenStudy (anonymous):

I honestly do not know. Haven't worked on it in a while. My interest change a lot

OpenStudy (anonymous):

This was incredibly strange though, given that completing the square was the only thing that was completed on the website and I just so happened to have broached this topic.

OpenStudy (anonymous):

It had quadratic formula also, but I might have deleted it or the link to it.

OpenStudy (anonymous):

A lot of time was just creating the infrastructure so that it would be easy to add in lessons. The infrastructure is there, but now it's hard to come up with concise lessons that don't feel awkward.

OpenStudy (anonymous):

I blame mathjax.

OpenStudy (anonymous):

It's not mathjax that is the problem.

OpenStudy (anonymous):

I will take your word for it. Time to ask more questions!

OpenStudy (anonymous):

I can pull up the code for that lesson, to show how simple it is.

OpenStudy (anonymous):

``` extends ../note-layout block vars - var topic = 'Completing the Square' block main +section('Synopsis') :markdown **Completing the square** is the process of converting a quadratic expression into its vertex form \\(a(x+h)^2+k\\) given its standard form \\(ax^2+bx+c\\). One way is to evaluate equations of \\(h\\) and \\(k\\) given the coefficients \\(a\\), \\(b\\), and \\(c\\). They are \\(h=\frac b{2a}\\) and \\(k=c-\left(\frac b{2a}\right)^2\\). .important :mathjax ax^2+bx+c = a\left(x+\frac b{2a}\right)^2+c-\left(\frac b{2a}\right)^2 :markdown In the case where \\(a=1\\) and \\(c\\) is ignored, we get \\(h=\frac b{2}\\) and \\(k=\left(\frac b{2}\right)^2\\). These equations are a bit easier to remember and can still be used even when \\(a\neq 1\\), so long as \\(a\\) is divided or factored out before completing the square. .important :mathjax x^2+bx = \left(x+\frac b{2}\right)^2-\left(\frac b{2}\right)^2 +section('Applications', {closed: true}) :markdown * Finding the roots of a quadratic equation. * Converting a quadratic equation vertex form. * Converting an equation of an ellipse, hyperbola, or parabola into its standard equation. ```

OpenStudy (anonymous):

What was the problem, then?